OpenStudy (anonymous):

If the average of three numbers is V. If one of the numbers is Z and another is Y, what is the remaining number? A. ZY - V B. Z/V - 3 - Y C. Z/3 - V - Y D. 3V- Z - Y E. V- Z - Y Answer is D. How so?

OpenStudy (anonymous):

You see average is calculated by the formula: Sum of quantities/No of quantities. So, (Z+Y+ ?)/3=V Thus, 3V= Z+Y+? 3V-Z-Y=? (? signifies the unknown number)

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

yw.